Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: In D configuration (all antennas moved as of the 26th; end of D is 27 December). All correlator hw resource issues resolved and ready (Q4 pair 15 is currently
reserved for testing purposes).
  • Key dates:
    • 26 Sep - 27 Dec: D configuration
  • Visitors:
    • Michiel Brentjens: 11 Jun - 11 Nov (Holography; characterizing polarized beams)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas: EA25 out.
    • Receivers:
      • EA24 for first L-band solar capable Rx.
      • Perley: RF Switch testing ongoing.
      • Calibration (Perley): Gain compression/expansion ongoing (Hayward).
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): On the sky acceptance document completed.
  • Computing (Robnett):
    • Notes:

Software status and updates (Butler)

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler); Note: Change for Mondays in D configuration - dedicated to science operations/planned commissioning tests.

  • Mon
    • 1000-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Tue:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday):
    • 0900-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: [Rupen/Carlson w/ Momjian/Sault] V stokes issues.Status
  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags). .Needs close-out testing (Michael-update activation time testing).
  • Correlator: [Clark/Carlson] Delay rocking across bands. (October)
    • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).

Testing This week/soon

  • Holography.
  • RSRO D configuration tests (principally data rate; recirc).
  • CM version update.
  • Subarrays: cleanup.
  • Stage 2 mixer testing to relax sub 128 MHz BW grids
  • Correlator: Rupen: S-band shadow testing
  • Spectral line: Roy/Dhawan: HI High resolution bandpass testing
  • Baseband edge/filter rolloff.
  • Sub-band dependent fshifts.
  • Array: *; Individual issue sleuthing.

Commissioning Groups

1000 -- 1150 MHz:  Pretty grim.  This is the zone of aeronautical
beacons (DMEs and radars).   The lines are all narrow, with an abundance
of 'holes'.
   1170 -- 1260 MHz:  Extremely strong and broad emission (GPS, and
others).
   1260 -- 1350 MHz:  Mostly empty, with a few (~8) narrow emissions
(radars, I think).
   1350 -- 1525 MHz:  Mostly empty.
   1525 -- 1610 MHz:  Zone of Death!  I doubt we'll ever be able to use
this region.
   1617 -- 1640 MHz:  Another deathly zone.
   1640 -- 1675:  Open
   1675 -- 1695:  Many RFI sources.
   1695 -- 1930:  Almost completely clear!
   1930 -- 1990:  Much RFI -- cellphones, I think (?)
   1990 -- 2180:  Mostly clear.
   2180 -- 2200:  Very strong RFI (geostationary satellites, I think).
   2200 -- 2310:  Mostly o.k., some strong RFI between 2225 -- 2240.
   2310 -- 2355:  Satellite radio. Completely hopeless!
   2355 -- 2500:  Mostly open, with a few (~8) narrow 'lines'.
   2500 -- 3580:  Mostly clear, but with strong RFI near:  2740, 2800
-- 2820, 3315, 3420 -- 3460 MHz.
   3580 -- 4200:  All of this span is heavily polluted -- I think this
is due to satellite downlinks for TV.
   4200 -- 6110:  Almost entirely open.  *** 5990 -- 6015 had strong
RFI in March 2010***
   6110 -- 6170:  Strong RFI throughout.  *** This span was entirely
open in March 2011***
   6170 -- 6360:  Open.  *** 6240 -- 6270 had strong RFI in March 2011***
   6360 -- 6420:  Strong RFI throughout.  *** This was entirely open in
March 2011***
   6612 -- 6618:  Strong RFI (which was also present in March 2011)
   6618 -- 6772:  Open.
   6772 -- 6777:   Strong RFI (also present in March 2011)
   6777 -- 7400:  Almost clear of RFI.
   7400 -- 7720:  Generally clear, but with scattered, generally
narrow, RFI.
   7720 -- 8000:  Clear of RFI.
    • 4-band (Owen):
      • Notes:
    • Adv: Low band (Owen):
      • Notes:
* Scheduling/Rapid Response (Claussen/McMullin): Link: https://safe.nrao.edu/wiki/bin/view/EVLA/RapidResponseObserving
      • Notes:

    • Adv: Astrometry (Mioduszewski):
      • Update:https://bugs.nrao.edu/browse/ECSV-173
  • Calibrator Models
    • Notes: Reduction underway; Ka/S coming very soon; others to follow.
  • HPC (Bhatnagar/Owen/Golap):
    • Notes: ~500 GB data set obtained. Issues noted when filling switched power.
  • Imaging (Ott/Golap):
    • Notes: Takahashi/Ott working on mosaic needs in PST/OPT. Document under review.
  • Algorithm (Bhatnagar, Owen):

RSRO/ECSO Reports

RSRO A configuration status: https://safe.nrao.edu/wiki/bin/view/EVLA/2011AConfiguration

Individual reports (updates)

Misc:

Demo Science Status

AOB

-- JosephMcMullin - 2011-09-28
Topic revision: r1 - 2011-09-28, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback